Circular No. 6669

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS G. E. Brueckner reports, for the SOHO-LASCO Consortium (which

consists of the Naval Research Laboratory, Laboratoire d’Astronomie Spatiale, Max-Planck-Institut für Aeronomie, and the University of Birmingham), the discovery of four more comets (cf. IAUC 6650, 6653) in images obtained by the LASCO white-light coronagraphs aboard the ESA/NASA Solar and Heliospheric Observatory (SOHO) spacecraft. The following approximate discovery positions are summarized from the more complete reductions by G. V. Williams on MPEC 1997-K10 to 1997-K13 from measurements by D. A. Biesecker and O. C. St. Cyr:

UT  α_{2000}  δ_{2000}
1996 Jan. 27.971  20^{h}56.^{m}4  −20^{o}01′  C/1996 B3
Dec. 11.056  17 15.8  −27 03  C/1996 X1
11.802  17 19.4  −27 11  C/1996 X2
1997 Jan. 26.179  20 54.2  −20 33  C/1997 B2

C/1996 B3 was discovered only recently using automated search software written by Biesecker. The others were discovered in real time by D. J. Lewis and Biesecker; S. T. Stezelberger; and Biesecker, respectively. The final analysis of the brightness of each object has not been completed, but preliminary results (based on comparisons with stars in the same coronagraphic field) by Biesecker and St. Cyr suggest that C/1996 X2 had m_{1} = 7.5 and the others m_{1} > 8.0, with variations evident for each object. Orbital computations by the undersigned on the same above-mentioned MPEC s suggest that all these SOHO comets are Kreutz sungrazers. with T = 1996 Jan. 28.71, Dec. 12.10, Dec. 12.88 and 1997 Jan. 27.08 UT, respectively. None of the four comets was detected receding from the sun, in spite of continuing observations in each case.

COMET 96P/MACHHOLZ 1 This comet was detected (accidentally by Stezelberger) during 1996

Oct. 13.4–14.7 UT near the edge of the SOHO LASCO C3 coronagraphic field. St. Cyr estimated m_{1} = 4 or 5, much brighter than the sungrazing comets reported above and on IAUC 6653. The positions reduced by Williams from St. Cyr’s sun-related measurements and published on MPC 29710 are within 2′ of this comet’s expected position and confirm the general correctness of the reduction procedure used for all of these comets.

K. J. Meech, O. R. Hainaut and J. M. Bauer, University of Hawaii, have reported observations of this comet with the 2.2-m reflector: 1997 Apr. 15.57 UT, m_{1} = 19.1; also observed on Apr. 17.6.

1997 May 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden